OpenStudy (quinton1012):

Forty-two is what percent of 70? A. 21% B. 60% C. 74% D. 167%

OpenStudy (maddy1251):

Take the number 70, and multiply it by all of the percentages. e.g. ; 70 x 0.21 = ?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

translate directly in to math

OpenStudy (misty1212):

42 is the number 42 "is" means = what percent is what you are looking for, say \(P\) "of" means times 70 is 70 \[42=P\times 70\] solve via \[P=42\div 70\] then turn your number in to a percent by moving the decimal two places right

OpenStudy (texaschic101):

is/of = %/100 is = 42 of = 70 % = x now sub 42/70 = x/100 70x = 4200 x = 4200/70 x = 60

OpenStudy (quinton1012):

thx guys